# HyperMOOC: Augmenting MOOC Videos with Concept-based Embedded Visualizations

### v1

@@ 3.3. Learning Content-level Design
Dimension I: Data Type. Existing work on augmented video points out that designers must consider striking a balance between explanatory and exploratory approaches  (Chen et al., 2023). Given that MOOC video learning is multi-stage (Mccarthy, 1996), learners typically start with explanatory content for basic concept acquisition and then apply their learning to exploratory tasks to gain a deeper understanding. Specifically, we notice that explanatory content usually has low semantics (e.g., text, figures, equations) that are directly embedded in the MOOC video and can be easily perceived, while exploratory data tend to have high semantics (e.g., concept relationships, concept importance) to provide learners with a deeper understanding of the course. Based on the differences in semantic levels between explanatory and exploratory data, we categorize the data in MOOC videos from low to high semantic levels:
@@
Element Level includes the most general representation of the concept, which can be obtained directly from the original video. Learners generally use this level of content to gain the most general knowledge of the course. Based on previous literature (Yadav et al., 2016; Chen and Wu, 2015; Chorianopoulos and Giannakos, 2013) and our review (Fig. 2-d), data at this level generally consists of basic elements describing the concept and auxiliary elements provided by the instructor. Specifically, these include: (1) Basic elements—such as text, figures, tables, equations, and code blocks; and (2) Auxiliary elements—such as teacher images, subtitles, tests, and examples. For instance, in a statistics course on MOOC, a textual definition of ”mean” displayed alongside an illustrative table constitutes Element Level data. Augmenting a video solely at this level is primarily for explanatory purposes.
@@
Event Level involves higher-level abstractions of concepts, which are generally difficult for learners to access directly from videos. Data at this level can further aid learners in understanding the patterns of knowledge dissemination and course organization. Providing additional presentation of this data is particularly effective for logic-driven and concept-intensive courses. Following the concept mapping method in MOOC videos (Zhou et al., 2024), we categorize this data into concept attributes and relationships. Attributes include duration, importance, and delivery style (i.e., whiteboard annotations, slide-based presentations, direct lecture delivery) and relationships consist of association, inclusion, and similarity. Augmenting data at this level makes videos both explanatory and exploratory.
@@
Conclusion Level presents course-level data, enhancing learners’ overall grasp of the course. This type of data explains the overall composition of the course and outlines the learning steps, such as the course’s type, structure, and important time nodes. Most current research based on learning and content analysis focuses on this level, using navigation menus (Zhou et al., 2024; Zhao et al., 2017) and course overview (Rahman et al., 2023; Schwab et al., 2016) to facilitate quick learning. Therefore, augmenting data (e.g., temporal structure, concept hierarchy) at this level is primarily for exploratory purposes, providing learners with additional information.
@@
Dimension II: Visual Effect. After summarizing all the data types in the videos an important question is how to augment these data through visualization. In the data video domain, data may be presented as graphical marks (Amini et al., 2017), and multiple graphical marks can be combined into information-rich frames to convey insights. Inspired by this, we consider augmenting data in videos as overlays of graphical marks. In addition, video effects are also widely used as a means of augmentation (Zhu-Tian et al., 2023; Chen et al., 2021, 2023), and different video effects can guide learners to redistribute their attention. We categorize these visual effects as graphical marks and video effects:
@@ 3.4. Learner-level Design
Dimension III: Behavior Level. In MOOC videos, the instructor’s teaching process tends to progress over time. However, we noticed that the process of learners’ learning does not strictly follow the time progression of the video. For example, learners control their pace by playing and pausing the video. Considering the general order of learners’ learning behaviors, we summarize previous works on learner behavior analysis (Ozan and Ozarslan, 2016; Shi et al., 2015; Sinha et al., 2014) to describe the main behaviors of learners:
Play denotes the first step of learning, where the learner follows the video through the instructor’s preset learning process. At this stage, learners tend to learn the most general, explanatory content.
Focused refers to the intermediate process of learning. Our user interviews showed that learners often moved their mouse over content, as being demonstrated by the instructor. Furthermore, informed by E1, we observed that learners tend to spend more time focusing on content they find difficult to understand, which aligns with Miller’s hypothesis of eye-movement cognitive engagement (Miller, 2015). Due to the limited availability and high cost of eye-tracking devices, we adopt a mouse-hover-based focus proxy to approximate gaze-based attention. This approach, primarily applied during the Play stage, allows us to infer areas of focus and trigger relevant augmentations. Learners typically deepen their understanding of explanatory content and seek out exploratory information at this stage.
Paused stage tends to promote the final understanding of concepts. Learners agreed that when the concepts they encountered were sufficiently difficult and complex, they often paused to explore them further. According to a previous clickstream analysis (Shi et al., 2015; Sinha et al., 2014), “seek” and “rate change” tend to be learners’ favorite actions. Moreover, the learning process recorded by one co-author shows that beyond videos, some learners preferred to use search engines such as Google, Wikipedia, or ChatGPT to gain a broader understanding of concepts, thus obtaining exploratory content for concept cognition.
Seek occurs mainly in the Paused stage, where learners drag the video timeline from one point to another to find supplementary content for concept explanation. However, learners often struggle to find the exact time point, and it usually takes several Seek operations.
@@
Dimension IV: Cognitive Level. The cognitive level consideration was mainly proposed by our experts and we categorize this level through cognitive load theory (Sweller et al., 1998). Experts agreed that the key strategy to enhance learning outcomes is to facilitate the reduction of various cognitive loads, freeing up more working memory for engaging with and understanding course content. In MOOC videos, factors such as content complexity, the organization of materials, and the ease of accessing auxiliary information affect understanding. Based on this, we characterize the main types of cognitive load that learners experience during the learning process:
@@ 4.1. Design Requirements
We further derived the design requirements of HyperMOOC based on the design space, previous research on augmented video and online learning (Chen et al., 2021; Zhou et al., 2024; Shi et al., 2015), and weekly meetings with two domain experts (E1 & E2, consistent with expert interviews) over the past 5 months. Notably, our design targets standard desktop environments because most current MOOC platforms are mainly accessed by online learners via computers. Ultimately, we grouped them into three categories: for explanatory purposes ( [Uncaptioned image] ), for exploratory purposes ( [Uncaptioned image] ) and for both explanatory and exploratory purposes ( [Uncaptioned image] ).
@@
    [Uncaptioned image] Provide an overview of concepts and enable real-time tracking. Existing analyses of learning content (Zhou et al., 2024; Schwab et al., 2016) often provide an overview of concepts as the initial step in learning. Additionally, augmented videos commonly employ Highlight effects for real-time augmentation (Chen et al., 2021, 2023). Therefore, E1 believed that enabling learners to understand the general course content and tracking the concept they are currently learning is the most basic augmentation that MOOC video learning should have.
@@
    [Uncaptioned image] Display elements that describe the concept. According to our design space, concepts have a variety of representations, explicitly displaying this information can help learners quickly access concept-related material, and the concept demonstration flow. This is crucial for promoting self-perceived learning.
@@
    [Uncaptioned image] Offer support for multi-stage learning behaviors. Given multiple learning behaviors, the augmented design should support different stage-related activities. For example, displaying the most basic, explanatory content while the video is playing, and providing exploratory content when the video pauses.
@@
Figure 3. The design of Play (a,b,c) and Focused (d) stage of HyperMOOC.
Figure 4. HyperMOOC interaction workflow includes 1) Play - Follow the instructor’s learning path. 2) Focused - Get concept explanation, and 3) Paused - Deepen the course understanding. Paused stage user interface, a) slides and course overview, b) video player, including (b1) concept relationships and (b2) demonstration materials, and c) video preview.
@@ 4.2. Usage Scenario
Get concept explanation. Currently, Alex is learning about the “histogram”, and he finds the concept of “percentages (relative frequencies)” somewhat challenging to grasp. To facilitate his understanding, the system allows him to hover the mouse over the text to retrieve the concept He finds the tool further enhances the Highlight effect, while all elements explaining the concept are displayed around the concept with icons (Fig. 3-d, R2). He can hover over these icons to view the original content of the elements and click to navigate. The same approach applies to other elements (e.g., figures, equations, etc.), with displayed content changing to the concepts contained within the elements. In this way, Alex can access concept-related material in real time to facilitate understanding.
@@ 4.4. Hyperlink-based Interaction Methods
Play. At this stage, HyperMOOC achieves the learner’s understanding of the current concept by [Uncaptioned image] Highlighting the concepts in the video frames (Fig. 3-b) and subtitles (Fig. 3-c). Learners can jump to important points in the course by clicking on important time nodes on the progress bar (Fig. 3-a), which is extremely beneficial for learners using HyperMOOC for review.
Focused. Given the limited prevalence and high complexity of eye-tracking devices, HyperMOOC currently implements a mouse-hover-based focus proxy as an alternative to eye-tracking. Specifically, when learners focus on a single element by hovering the mouse over it for more than 3 seconds, the tool prompts further augmentation of that element. It allows the system to identify the user’s focus and provide additional explanatory information or highlighting around the concept they are attempting to understand.
Paused. Learners Pause the video and switch to HyperMOOC’s full mode by clicking on elements within the video. The full mode globally supports two Keyboard-mouse interaction modes: hovering and clicking. Learners can hover over concepts or other elements to display their original content or click to navigate. When the hovered element is a concept, the content in the Demonstration area changes to the explanation flow of the selected concept (Fig. 4-b2), facilitating a continuous understanding of the knowledge. Additionally, learners can select the currently focused concept using the video preview (Fig. 4-c).
@@ 5.1. User Study
Participants and Apparatus. We recruited 36 participants (20 males and 16 females) via university mailing lists and forums. Participants were evenly distributed across three academic majors—Mathematics, Education, and Art Design—with 12 from each. Within each major, four participants were randomly assigned to one of three learning modes. Mathematics participants were highly familiar with the content, Education participants had partial background knowledge, and Art Design participants were encountering this material for the first time. The average age was 23.9 years (SD = 2.57). The experiment was conducted in a lab using a 24-inch monitor (1920×1080 resolution). Each session lasted approximately 70 minutes, and participants were compensated \$6.
@@
Figure 8. Learning test results of three participant groups across three learning modes in two tasks (the error bars indicate the standard deviation (SD) of the mean scores across participants). Task 1 – RAW vs. AUG, Task 2 - AUG vs. FULL.
@@
Learning Outcomes. We analyzed quiz scores from the three participant groups across the three learning modes over two tasks. The results showed significant differences among groups and modes (Fig. 8). The FULL mode of HyperMOOC was most effective in enhancing learning outcomes, especially for Education majors. This can be attributed to their stronger theoretical background and greater familiarity with pedagogical content, which likely made them more receptive to the structured, explanation-rich format of the FULL mode. Art Design students also showed significant improvements in the AUG and FULL modes compared to the RAW mode. Although their overall scores remained lower than those of the other groups, their learning gains suggest that the enhanced visual and interactive elements in these modes better aligned with their learning preferences.
@@
Overall User Experience. The results of the SUS (Fig. 9 left) indicate that the majority of participants felt at ease using HyperMOOC and showed a preference for utilizing it in their MOOC video learning. Fig. 9 right shows the ranking of the three modes in terms of reducing cognitive load. Fifteen participants (62.5%) believed that FULL mode could significantly reduce cognitive load, followed by AUG mode with nine participants (41.7%). The remaining nine participants (37.2%) did not list FULL as their top choice, mainly expressing concerns that the embedded visualizations provided by HyperMOOC might increase learners’ cognitive load.
User experiences varied significantly based on academic backgrounds. HyperMOOC was effective for students with relevant foundational knowledge but challenging for students lacking prior knowledge. One Art Design student reported significant difficulties using the system. Most users adapted well to embedded visualization features, but a minority, mainly from Art Design, expressed concerns about advanced features potentially increasing cognitive load. These findings emphasize the importance of considering user background diversity in online learning system design, especially for interdisciplinary content. They provide direction for optimizing HyperMOOC and highlight the need to support learners with different knowledge backgrounds through flexible knowledge presentations or personalized learning pathways.

### v2

@@ 3.3. Learning Content-level Design
Dimension I: Data Type. The derivation of this dimension triangulates three sources: literature on video augmentation establishing the need to balance explanatory and exploratory approaches (Chen et al., 2023), our observation that 79.7% of analyzed papers distinguish between content-level and relationship-level data, and user interviews where 8 out of 10 participants explicitly differentiated “understanding what’s on the slide” versus “understanding how concepts connect.” Given that MOOC learning is multi-stage (Mccarthy, 1996), learners first acquire basic concepts before applying knowledge in deeper exploration. We thus identified a semantic hierarchy: explanatory content exhibits low semantic complexity (text, figures, equations directly embedded), while exploratory data carry high semantic complexity (concept relationships, importance indicators requiring deeper cognitive processing). Expert E1 confirmed: “students first need to see what’s there before understanding how it all fits together.” This semantic stratification provides the foundation for categorizing MOOC video data from low to high levels:
@@
[Uncaptioned image]Element Level includes the most general representation of the concept, which can be obtained directly from the original video. Learners generally use this level of content to gain the most general knowledge of the course. Based on previous literature (Yadav et al., 2016; Chen and Wu, 2015; Chorianopoulos and Giannakos, 2013) and our review (Fig. 2-d), data at this level generally consists of basic elements describing the concept and auxiliary elements provided by the instructor. Specifically, these include: (1) Basic elements—such as text, figures, tables, equations, and code blocks; and (2) Auxiliary elements—such as teacher images, subtitles, tests, and examples. For instance, in a statistics course on MOOC, a textual definition of ”mean” displayed alongside an illustrative table constitutes Element Level data. Augmenting a video solely at this level is primarily for explanatory purposes.
@@
[Uncaptioned image]Event Level involves higher-level abstractions of concepts, which are generally difficult for learners to access directly from videos. The extraction of key concepts and the establishment of their structured relationships to enhance learners’ understanding has also garnered significant attention from learners(Zhang et al., 2019; Zhao et al., 2017). Donald(Donald, 1983) categorized the relationships between concepts into three types: associative, functional, and structural, which are further mapped to Sequence, Association, Similarity, and Inclusion to better express the relationships between concepts in MOOC videos. Providing additional presentation of this data is particularly effective for logic-driven and concept-intensive courses (Zhou et al., 2024). Thus, we categorize this data into concept attributes and relationships. Attributes include duration, importance, and delivery style (i.e., whiteboard annotations, slide-based presentations, direct lecture delivery) and relationships consist of association, inclusion, and similarity. Augmenting data at this level makes videos both explanatory and exploratory.
@@
[Uncaptioned image]Conclusion Level presents course-level data, enhancing learners’ overall grasp of the course. This type of data explains the overall composition of the course and outlines the learning steps, such as the course’s type, structure, and important time nodes. Most current research based on learning and content analysis focuses on this level, using navigation menus (Zhou et al., 2024; Zhao et al., 2017) and course overview (Rahman et al., 2023; Schwab et al., 2016) to facilitate quick learning. Therefore, augmenting data (e.g., temporal structure, concept hierarchy) at this level is primarily for exploratory purposes, providing learners with additional information.
@@
Dimension II: Visual Effect. Having established what to augment, this dimension addresses how. In the data video domain, data may be presented as graphical marks (Amini et al., 2017), combined into information-rich frames to convey insights. User interviews validated this framework as participants naturally distinguished “seeing additional information” versus “being guided to look at something,” while E2 characterized it as “showing versus highlighting.” Drawing from visualization grammar (Amini et al., 2017) and video augmentation literature (Zhu-Tian et al., 2023; Chen et al., 2021, 2023), we categorize visual effects as: (1) graphical marks providing persistent information overlays, and (2) video effects dynamically redirecting attention.
@@ 3.4. Learner-level Design
Dimension III: Behavior Level. This dimension emerged from a critical observation across our data: while instructors design MOOC videos with temporal linearity, learner behavior is non-linear and interactive. All 10 interview participants demonstrated varied engagement patterns—pausing, rewinding, and skipping—validated by literature on learner behavior (Ozan and Ozarslan, 2016; Shi et al., 2015; Sinha et al., 2014) documenting consistent interaction patterns beyond sequential viewing. Expert confirmation strengthened this finding: E1 noted “students rarely just watch straight through—they pause, they jump around,” while E2’s analytics showed “video interaction patterns strongly predict learning outcomes.” Synthesizing literature, interview observations, and expert validation, we derived three fundamental behavioral states capturing how learners actively control their learning process from passive consumption to active exploration:
[Uncaptioned image]Play denotes the first step of learning, where the learner follows the video through the instructor’s preset learning process. At this stage, learners tend to learn the most general, explanatory content.
[Uncaptioned image]Focused refers to the intermediate process of learning. Our user interviews showed that learners often moved their mouse over content being demonstrated by the instructor. Informed by E1, we observed that learners tend to spend more time focusing on difficult content, aligning with Miller’s hypothesis of eye-movement cognitive engagement (Miller, 2015). While eye-tracking provides more precise attention measures, we adopt a mouse-hover-based focus proxy considering ecological validity: most MOOC learning occurs on personal devices without eye-trackers, and mouse cursor movements correlate with visual attention (Huang et al., 2011). This approach allows us to infer areas of focus and trigger relevant augmentations during the Focused stage. Learners typically deepen their understanding of explanatory content and seek out exploratory information at this stage.
[Uncaptioned image]Paused stage tends to promote the final understanding of concepts. Learners agreed that when the concepts they encountered were sufficiently difficult and complex, they often paused to explore them further. According to a previous clickstream analysis (Shi et al., 2015; Sinha et al., 2014), “seek” and “rate change” tend to be learners’ favorite actions. Moreover, the learning process recorded by one co-author shows that beyond videos, some learners preferred to use search engines such as Google, Wikipedia, or ChatGPT to gain a broader understanding of concepts, thus obtaining exploratory content for concept cognition.
[Uncaptioned image]Seek occurs mainly in the Paused stage, where learners drag the video timeline from one point to another to find supplementary content for concept explanation. However, learners often struggle to find the exact time point, and it usually takes several Seek operations.
@@
Dimension IV: Cognitive Level. While prior dimensions emerged from literature and user data, this was explicitly proposed by both domain experts as a critical perspective. During consultations, E1 and E2 emphasized cognitive load management—E1 noting “adding more information can sometimes make things worse,” and E2 citing research linking cognitive load to MOOC completion rates. This directed us to cognitive load theory (Sweller et al., 1998). We categorize loads by their sources and instructional implications, distinguishing between unavoidable content difficulty, processing overhead from poor presentation, and productive cognitive effort for schema construction:
@@ 4.1. Design Requirements
We further derived the design requirements of HyperMOOC based on the design space, previous research on augmented video and online learning (Chen et al., 2021; Zhou et al., 2024; Shi et al., 2015), and weekly meetings with two domain experts (E1 & E2, consistent with expert interviews) over the past 5 months. Notably, the experts’ insights were grounded in systematic observations of learner behaviors from discussion forums and post-class feedback throughout their teaching experience. Besides, our design targets standard desktop environments because most current MOOC platforms are mainly accessed by online learners via computers. Ultimately, we grouped them into three categories: for explanatory purposes ( [Uncaptioned image] ), for exploratory purposes ( [Uncaptioned image] ) and for both explanatory and exploratory purposes ( [Uncaptioned image] ).
@@
    [Uncaptioned image] Provide an overview of concepts and enable real-time tracking. Existing analyses of learning content (Zhou et al., 2024; Schwab et al., 2016) provide concept overviews as the initial learning step, and augmented videos commonly employ Highlight effects for real-time augmentation (Chen et al., 2021, 2023). E1 observed that learners frequently posted forum questions like “where are we in the course?” and post-class feedback confirmed that nearly 70% of learners reported difficulty tracking course structure. Thus, enabling learners to understand general course content and track their current concept is the most fundamental MOOC video augmentation.
@@
    [Uncaptioned image] Display elements that describe the concept. According to our design space, concepts have a variety of representations. E1 noted learners’ common forum complaints about difficulty locating “which slide had that example,” and mentioned frequently “need pausing to find specific content.” Explicitly displaying this information can help learners quickly access concept-related material and the concept demonstration flow, which is crucial for promoting self-perceived learning.
@@
    [Uncaptioned image] Offer support for multi-stage learning behaviors. Given multiple learning behaviors observed by both experts, the augmented design should support different stage-related activities through seamlessly embedded visualizations. Specifically, the system should provide real-time, low-interruption augmentation during video playback (Play and Focused stages) and enable on-demand exploration when learners pause, all within the video canvas to maintain spatial continuity and minimize attention shifts.
@@
Figure 3. The design of Play (a,b,c) and Focused (d) stage of HyperMOOC, Alex is currently viewing the original content of concept “Percentages” in “Preparation” stage.
Figure 4. HyperMOOC interaction workflow includes 1) Play - Follow the instructor’s learning path. 2) Focused - Get concept explanation, and 3) Paused - Deepen the course understanding. Paused stage user interface, a) slides and course overview, b) video player, including (b1) concept relationships and (b2) demonstration materials, and c) video preview. Alex is currently delving into the concept of “Percentages” and how it relates to others.
@@ 4.2. Usage Scenario
Get concept explanation. Currently, Alex is learning about the “histogram”, and he finds the concept of “percentages (relative frequencies)” somewhat challenging to grasp. To facilitate his understanding, the system allows him to hover the mouse over the text to retrieve the concept. He finds the tool further enhances the Highlight effect, while all elements explaining the concept are displayed around the concept with icons (Fig. 3-d, R2). He can hover over these icons to view the original content of the elements and click to navigate. The same approach applies to other elements (e.g., figures, equations, etc.), with displayed content changing to the concepts contained within the elements. In this way, Alex can access concept-related material in real time to facilitate understanding.
@@ 4.4. Hyperlink-based Interaction Methods
[Uncaptioned image]Play (Real-time Augmentation). At this stage, HyperMOOC provides synchronized visual cues that update in real-time during video playback, [Uncaptioned image] Highlighting the concepts in the video frames (Fig. 3-b) and subtitles (Fig. 3-c). Learners can jump to important points in the course by clicking on important time nodes on the progress bar (Fig. 3-a), which is extremely beneficial for learners using HyperMOOC for review.
[Uncaptioned image]Focused (Real-time Detail-on-Demand). Given the limited prevalence and high complexity of eye-tracking devices, HyperMOOC currently implements a mouse-hover-based focus proxy as an alternative to eye-tracking. Specifically, when learners focus on a single element by hovering the mouse over it for more than 3 seconds, the tool prompts further augmentation of that element. It allows the system to identify the user’s focus and provide additional explanatory information or highlighting around the concept they are attempting to understand.
[Uncaptioned image]Paused (On-demand Exploration). When learners Pause the video, HyperMOOC seamlessly transitions to its full exploratory mode, expanding the embedded visualizations to reveal comprehensive concept relationships and cognitive paths. The full mode globally supports two Keyboard-mouse interaction modes: hovering and clicking. Learners can hover over concepts or other elements to display their original content or click to navigate. When the hovered element is a concept, the content in the Demonstration area changes to the explanation flow of the selected concept (Fig. 4-b2), facilitating a continuous understanding of the knowledge. Additionally, learners can select the currently focused concept using the video preview (Fig. 4-c).
@@ 5.1. User Study
Participants and Apparatus. We recruited 36 participants (20 males, 16 females) via university mailing lists and forums. This sample size was determined to provide sufficient statistical power for our between-subjects design to detect significant differences in learning outcomes and cognitive load, balancing quantitative statistical requirements with the depth of qualitative feedback required for the study. This scale aligns with contextual practices in similar augmented video research (e.g., Chen et al. (Chen et al., 2023) used 16 participants), ensuring comparable rigor. Participants were evenly distributed across three academic majors—Mathematics, Education, and Art Design—with 12 from each. Within each major, four participants were randomly assigned to one of three learning modes. Mathematics participants were highly familiar with the content, Education participants had partial background knowledge, and Art Design participants were encountering the material for the first time. The average age was 23.9 years (SD = 2.57). The experiment was conducted in a lab using a 24-inch monitor (1920×1080 resolution). Each session lasted approximately 70 minutes, and participants were compensated \$6.
@@
Figure 8. Learning test results of three participant groups across three learning modes in two tasks (the error bars indicate the standard deviation (SD) of the mean scores across participants).
@@
Learning Outcomes. One-way ANOVA revealed significant mode effects for Task 1 (F(2, 33) = 7.08, p = 0.003, $\eta^{2}$ = 0.30) and Task 2 (F(2, 33) = 5.90, p = 0.006, $\eta^{2}$ = 0.26) . Bonferroni-corrected comparisons (APPENDIX F-Table 1) showed FULL mode significantly outperformed RAW in both tasks (Task 1: p=0.010, d=1.34; Task 2: p=0.013, d=1.30), with significance markers shown in Fig. 8. FULL also exceeded AUG in Task 1 (p=0.012, d=1.31). Exploratory analysis revealed Education majors showed greatest improvement with FULL mode (moderate prior knowledge enabled them to leverage structured visualizations), followed by Art Design (benefited from enhanced visual support despite lower baseline) and Mathematics majors (smaller gains due to existing strong knowledge).
Overall User Experience. The results of the SUS (Fig. 9 left) indicate that the majority of participants felt at ease using HyperMOOC and showed a preference for utilizing it in their MOOC video learning. Fig. 9 right shows the ranking of the three modes in terms of reducing cognitive load. Twenty-three participants (63.9%) believed that FULL mode could significantly reduce cognitive load, followed by AUG mode with fourteen participants (38.9%). The remaining thirteen participants (36.1%) did not list FULL as their top choice, mainly expressing concerns that the embedded visualizations provided by HyperMOOC might increase learners’ cognitive load.
User experiences varied significantly based on academic backgrounds. HyperMOOC was effective for students with relevant foundational knowledge but challenging for students lacking prior knowledge. One Art Design student reported significant difficulties using the system. Most users adapted well to embedded visualization features, but a minority, mainly from Art Design, expressed concerns about advanced features potentially increasing cognitive load.
To understand these differences, we analyzed NASA-TLX scores by dimension and participant major. FULL mode (M=33.8, SD=6.7) was rated significantly lower than RAW (M=54.9, SD=6.3), indicating reduced cognitive effort in comprehension. However, this benefit varied by background: Mathematics majors rated FULL mode lowest on Mental Demand (M=29.2), while Art Design majors rated it highest (M=46.7), a 17.5-point difference (p ¡ 0.001). Post-hoc interviews revealed that Art Design participants found the radial visualization initially “overwhelming” (P21), whereas Mathematics participants appreciated being able to “see the whole structure at once” (P5). This suggests that domain familiarity moderates the effectiveness of complex embedded visualizations.
These findings emphasize the importance of considering user background diversity in online learning system design, especially for interdisciplinary content. They provide direction for optimizing HyperMOOC and highlight the need to support learners with different knowledge backgrounds through flexible knowledge presentations or personalized learning pathways.